Consider the following  information:  

Marginal propensity to consume = 0,9

Investment = R180 million

Autonomous consumption = R105 million. The equilibrium level of income is:-

Question 2 options:


1800 million.


2800 million.


0,9Y.


2850 million


1
Expert's answer
2021-12-07T18:52:18-0500

"Y=Ca+MPC\times Y+I"


"Y=105+0.8\times Y+180"


"Y-0.8\times Y=285"


0.2Y=285


Y=1425